OpenStudy (anonymous):

‎1) Integral of 1/ ( x^2 - x ) dx 2) x / (x^2+1) dx = (y^2 + 1) / (y+1) dy

OpenStudy (anonymous):

We'll decompose the fraction 1/ ( x^2 - x ) into elementary fractions Int dx/ ( x^2 - x ) = -Int dx/x + Int dx/(x-1) Int dx/ ( x^2 - x ) = -ln|x| + ln|x-1| + C Int dx/ ( x^2 - x ) = ln|(x-1)/x| + C

OpenStudy (anonymous):

Thank you.